OpenStudy (anonymous):

find the quadratic equation that fits the following data points {(5,246),(-1,0)(4,160)}

OpenStudy (amistre64):

i dont understand the notation

OpenStudy (amistre64):

is that just 3 points?

OpenStudy (amistre64):

i got something I made just for this lol

OpenStudy (anonymous):

{(5,246),(-1,0),(4,160)}

OpenStudy (amistre64):

we plug the points in and it spits a quad at us

OpenStudy (anonymous):

ok but what do i plug in

OpenStudy (amistre64):

9x^2 + 5x -4

OpenStudy (amistre64):

the data points

OpenStudy (anonymous):

where did u get 9x^2+5x-4

OpenStudy (amistre64):

well, I programed it to perform all that tedious calculations that I would do by hand from the data points and then tell me what I would have come to as a result

OpenStudy (anonymous):

i see u did the work for me

OpenStudy (amistre64):

its alot of solving for a system of equations in 3 variables

OpenStudy (anonymous):

thats what i was trying to do it sucks i always get stuck on whether to multiply (4) and (5)

OpenStudy (anonymous):

so to write it in standard y=a^2+bx+c would be y=9x^2+5x-4

OpenStudy (amistre64):

the equations are: 5^2a +5b +c =246 -1^2a -1b +c = 0 4^2a +4b +c =160

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

great

OpenStudy (anonymous):

as usuall your a genious